# Remote Cooling of Spin-ensembles Coupled to a Spin-mechanical Hybrid Interface

### v1

@@ Abstract
We present here a protocol for the ground-state cooling of a hybrid quantum system consisting of a spin ensemble, an oscillator, and a single probe spin. In the weak dispersive coupling limit between the spins and the oscillator, a back-action effect from frequent measurements of the probe spin allows exponential cooling of the hybrid system. We identify the parameter regimes necessary to cool the ensemble, the oscillator, or both to their thermal ground states. Our results contribute to the development of a practical solution for cooling/polarizing large spin ensembles. More importantly, we show that even with weak coupling, significant changes in the dynamics of such ensemble-oscillator hybrid systems can be achieved by manipulating a single spin. This result provides a relatively simple means of tuning the dynamics of a hybrid system, which might facilitate advances in many quantum technology applications such as preparing entangled states in macroscopic objects.
@@ I Introduction
Quantum technology holds great potential to revolutionize areas such as information processing [1, 2, 3], secure communication [4, 5], sensing [6, 7], and fundamental testing [8]. However, translating this potential into real-world applications remains a significant challenge due to the inherent limitations of current experimental platforms [9, 10]. For example, nitrogen-vacancy (NV) centers in diamond [11] have an extended coherence time [12, 13] and an optical interface that enables long-range entanglement [14, 15], but their limited control accuracy [16] hinders advanced applications such as fault-tolerant quantum computation [17, 18]. Therefore, it is essential to explore the hybrid system approach that fully exploits the unique capabilities of different systems [19, 20, 21, 22]. Following this idea, large NV ensembles have been used exclusively as memory for superconducting qubits [23, 24], known for their superior control fidelity [1], reducing the need for high-precision entangling gates over NV centers required for quantum computation. The rich dynamics of hybrid quantum systems also opens up significant prospects for sensing [25, 22], simulation [21, 26] and machine learning [27, 28, 29].
Establishing high-fidelity entanglement between remote spin qubits through a bus is a well-studied paradigm [30, 22]. The idea is to transduce their common coupling to macroscopic objects (bus) into indirect interactions that lead to their entanglement [30, 31]. Nano-mechanical oscillators have been playing a pivotal role in creating universal quantum transducers for spin–spin spin-photon and phonon-photon interactions [32, 33]. Other transducers in the form of magnons [34], superconducting qubits [35, 36], optical cavities [25], optomechanical cavities [37], and superconducting resonator [38] have also been considered. There is also a proposal of using a three-level spin as the transducer to mediate entanglement between two oscillators [39].
@@
Transducers have primarily been proposed to create indirect interactions between the initially non-interacting systems, effectively implementing entanglement gates. between remote qubits [37, 33, 34]. Nevertheless, there are few transducer protocols that focus on heat extraction from quantum systems to achieve ground state cooling or quantum state preparation. This aspect is crucial, especially for large spin ensembles, since conventional ensemble cooling or polarization techniques often rely on resonant energy exchange methods, such as the Hartman-Hahn transfer [40, 41].
Polarization based on energy transfer typically exploit the interaction between a single spin and all the ensemble spins. This single spin can be repeatedly polarized, thus realizing a continuous removal of entropy from the ensemble, leaving it finally polarized. However, when dealing with identical spins (with similar g-factors), the risk of the ensemble being trapped in dark states due to Hamiltonian symmetries is significantly high [41, 42], which prevents full polarization. This dark state problem can be overcome by performing projective measurements on a single qubit interacting with the target ensemble and using its back action to extract heat. Such quantum state projections of a single spin not only allow efficient heat extraction from an ensemble, but can also work in a non-trivial situation where the original system dynamics does not provide any energy exchange [43, 44].
In this work, we will consider a tripartite hybrid system consisting of a spin ensemble, an oscillator, and a single spin (probe). Since we envision the oscillator to be a macroscopic object such as a nano-mechanical oscillator, the probe spin and the ensemble can be remotely entangled, as schematically shown in Fig. 1. In the limit where the oscillator is dispersively coupled (off-resonant) to the probe and ensemble spins, no energy exchange is possible from the natural Hamiltonian dynamics. However, by exploiting the measurement back action from the frequently measured probe spin, it is possible to cool/polarize the spin ensemble from its initial fully mixed thermal states to the ground states.
Our cooling protocol is based on indirect coupling between the probe spin and the ensemble that is mediated by the oscillator. It has already been demonstrated that multiple spins can be coupled to a nano-mechanical oscillator [43, 44], which would be much easier than achieving strong coupling between a single spin and all target spins in a large ensemble directly [24]. Moreover, since only weak dispersive coupling between the spins and the oscillator is required, our protocol can be straightforwardly implemented on a large variety of physical systems and is anticipated to be more feasible for cooling down larger electron-spin ensembles. Furthermore, given the wide operating temperature range of contemporary hybrid systems, from $4$K to $300$K, the need for cooling down individual components is paramount [45].
@@
The paper is organized as follows: In Sec. II, we outline the Hamiltonian of the system and discuss possible experimental implementations of our proposal. In Sec. III, we analyze the dynamics of the hybrid system and describe how periodic dynamical decoupling (DD) pulses on the probe spin can modify the dynamics of other parts in the system. In Sec. V, we study the parameters that allow cooling of the spin ensemble, the oscillator, or both. Finally, Sec. VI provides a summary and discusses possible directions for future research.
II SETUP
This work focuses on a tripartite quantum system consisting of an ancilla spin (probe), a spin ensemble, and an oscillator. This hybrid system can be realized as shown in Fig. 1 through a one-sided clamped cantilever embedding a single NV center positioned such that it experiences maximal strain coupling. To the edge of the cantilever is attached a magnetic tip that produces a field gradient $\nabla B$. This field is then coupled to a spin ensemble on an external substrate. The spin ensemble and the probe spin are both coupled to the mechanical motion through the field gradients [33, 22] or longitudinal strain [46, 47, 48].
@@ II SETUP
For simplicity, we assume (though not necessary) that the ensemble spins interact with the common mechanical mode uniformly with a coupling strength denoted as $g$. In addition, the probe spin is assumed to have a distinct coupling strength with the oscillator, denoted as $g_{0}$.
The hybrid system’s Hamiltonian is then formulated as follows:
@@
Figure 1: (a) Schematic illustration of the control procedure detailed in the paper. (a) A clamped nano/micro-mechanical cantilever hosting single defect spin, that is coupled to its motion through strain and is manipulated by laser and RF fields. A magnetic tip attached at the edge of the cantilever produces a gradient field $\nabla\vec{B}$ on a nearby spin ensemble allowing for its coupling to the mechanical motion of the cantilever. (b) Schematic illustration of the coupling and control mechanism for the above described spin-mechanical system. While the probe spin (green) is reset (to state |$\ket{0}$) after every projective measurement, the oscillator ($\rho_{o}$) and ensemble states ($\rho_{S}$) continue to evolve conditioned on the measurement of the probe.
@@
The single probe spin is envisioned to be solid state spins in diamond such as NV center electronic spins, due to good spin coherence properties, high degree spin control and a well-resolved optical spectrum at low temperatures [49, 17, 16]. The optical interface allows optical excitation to various levels that can either initialize or readout the spin state with fidelities exceeding $98\%$ [50]. In addition, these spins can also (weakly) couple to external spins, photons [43, 51] and phonons [44, 37].
@@
Note that the NV electronic spin-1 can be operated as a qubit with its $m_{s}=0,-1$ sub-levels by applying a 400 Gauss external magnetic field aligned with its symmetry axis to separate the $m_{s}=\pm 1$ levels [17]. Such an arrangement negates the need to consider transverse spin-strain coupling due to its direct interaction between the $m_{s}=\pm 1$ levels [30, 47, 48], so we can safely ignore the $m_{s}=+1$ sublevel in this work. What remains is the longitudinal strain coupling acting on the qubit levels. This number can be up to 10 kHz for NV centers [52], and higher for other defect centers [53].
@@
The spin ensemble to be cooled can be chosen with more flexibility, as stringent control over these spins is not a prerequisite for the current study. Potential candidates for the spin ensemble include distant spins on other coupled cantilevers, spins at the clamped edge that are strain-coupled to the mechanical mode, or other surface electron spins. Even more complex would be the situation where the spins coupled to the mechanical mode are not the spins attached to the NV centers in diamond but are dark (electronic) spins on the surface [54] of the NMO, or other ferromagnetic particles [55]. To harness those spins for spin-mechanics one should rely on indirect polarization techniques as they cannot be directly polarized optically. Note taht, as schematically shown in Fig. 1(a), the interaction between the ensemble spins and the mechanical vibrational mode is generated by a magnetic tip attached to the cantilever with high field gradients. Here the coupling strength can be much higher [56], than the strain coupling between probe and the oscillator.
III System dynamics
@@
Since the ensemble spins form a finite-dimensional Hilbert space, the system evolution operators can be written in the form [57, 58]:
@@
where $\mathcal{T}$ the time ordering operator, and $\mathrm{I}_{\scriptscriptstyle k}$ projects the spin ensemble into the subspace where $k$ spins are pointing up. For example, when $N=3$ and $k=2$, the projector is written as
@@
The term $h_{\scriptscriptstyle k,\pm}(t)$ is hence the oscillator Hamiltonian conditioning on the states if the spins. It is expressed as
$h_{\scriptscriptstyle k,\pm}(t)=\frac{g(k-N/2)\pm g_{0}}{2}\left(\hat{a}e^{-i\omega t}+\hat{a}^{\dagger}e^{i\omega t}\right).$ (5)
@@
where the functions $\alpha(t)$, $\beta(t)$ and $\theta_{\scriptscriptstyle k}$ are given by
$\begin{split}\alpha(t)&=1-e^{i\omega t},\quad\beta(t)=\frac{\omega t-sin(\omega t)}{\omega t},\\ \theta_{\scriptscriptstyle k}&=-\beta(t)\left(g(k-N/2)\pm g_{0}\right)^{2}/4\omega.\end{split}$ (7)
IV Probe spin control
In addition to the natural dynamics induced by the spin-oscillator interaction given in Eq. (1), we also apply a periodic sequence of $\pi$ pulses on the probe spin. These pulses significantly alter the dynamics of the system by embedding a time-periodic function $f(t)$ into the Hamiltonian, described as follows:
@@ IV Probe spin control
where $j=0,1,2,\cdots,n$ and $\tau$ denotes the interval between successive pulses. By replacing $g_{0}$ with $f(t)g_{0}$ in $h_{\scriptscriptstyle k,\pm}(t)$,the conditional evolution operators in Eq (6) are now written as
$\begin{split}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)=&\exp\Big{[}\frac{g(k-N/2)}{2\omega}\left(\alpha(t)a^{\dagger}-\alpha^{*}(t)a\right)\\ &\;\pm\frac{f(t)g_{0}}{2\omega}\left(F(\epsilon,t)a^{\dagger}-F^{*}(\epsilon,t)a\right)\Big{]}\exp\left[i\theta_{\scriptscriptstyle k}(t)\right],\end{split}$ (9)
@@
$\begin{split}F(t)&=i\omega\ \sum_{k=0}^{n}\int\limits_{k\tau}^{(k+1)\tau}\mathrm{d}t\;(-1)^{k}e^{-i\omega t}\\ &=-\sum_{k=0}^{n}(-1)^{k}\left[e^{-i(k+1)\omega\tau}-e^{-ik\omega\tau}\right]\\ &=\left[1-(-1)^{n+1}e^{-i(n+1)\omega\tau}+2\sum_{k=1}^{n}(-1)^{k}e^{-ik\omega\tau}\right].\end{split}$ (10)
This filter function can be expressed in a well-known form, centered at $\tau=\pi/\omega$ and with its width decreasing approximately as $1/n^{2}$ [58, 57, 59]:
$F^{2}(\tau,n,\omega)=4\tan^{2}\left(\frac{\omega\tau}{2n+2}\right)\cos^{2}\left(\frac{\omega(n+1)\tau}{2}\right),$ (11)
The asymptotic behavior of the filter function is given by
$\lim\limits_{n\rightarrow\infty}F(t)=2(n+1)\delta(\omega-\frac{\pi}{\tau}).$ (12)
In practice, to ensure only the mode at the particular frequency $\pi/\tau$ couples to the probe spin, the number of pulses on the probe should be $n>\mathcal{O}(\omega/g_{0}$). Other phonon modes are effectively decoupled [58], which supports our initial assumption of a single oscillator mode in Eq. (1).
Furthermore, our analysis focuses on the scenario where the inversion pulses resonate with the oscillator frequency. To achieve resonance, we set the interval between successive inversion pulses as
$\tau=\frac{\pi}{\omega-\epsilon},\text{ where }\frac{\epsilon}{\omega}\ll 1,$ (13)
where the detuning parameter $\epsilon$ is intentionally kept significantly smaller than the oscillator frequency $\omega$ to facilitate near-resonant conditions. The filter function is now simplified to

### v2

# Remote Cooling of Spin-ensembles through a Spin-mechanical Hybrid Interface

@@ Abstract
We present a protocol for the ground-state cooling of a tripartite hybrid quantum system, in which a macroscopic oscillator acts as a mediator between a single probe spin and a remote spin ensemble. In the presence of weak dispersive coupling between the spins and the oscillator, cooling of the oscillator and the ensemble spins can be achieved by exploiting the feedback from frequent measurements of the single probe spin. We explore the parameter regimes necessary to cool the ensemble, the oscillator, or both to their thermal ground states. This novel cooling protocol shows that, even with only weak dispersive coupling, energy transfer-like effects can be obtained by simply manipulating the probe spin. These results not only contribute to the development of a practical solution for cooling/polarizing large spin ensembles, but also provide a relatively simple means of tuning the dynamics of a hybrid system. The proposed cooling protocol thus has broader implications for advancing various quantum technology applications, such as macroscopic quantum state generation and remote sensing.
@@ I Introduction
Translating the potential of quantum technology into real-world applications remains a significant challenge due to the inherent limitations of current experimental platforms [1, 2]. One promising avenue is to explore the hybrid system approach, which promises to fully exploit the unique capabilities of different systems [3, 4, 5, 6]. In particular, high-fidelity entanglement between different systems can be established by exploiting their common coupling to a transducer, thereby generating indirect interactions [7, 6]. Transducers in the forms of nanomechanical oscillators (NMOs) [8, 9], magnons [10, 11], superconducting qubits [12, 13], superconducting resonators [14], optical cavities [15], and optomechanical cavities [16] have been considered to mediate spin–spin, spin-photon, and photon-phonon interactions.
@@
However, it has been mainly proposed to use transducers to implement entanglement gates between remote qubits [16, 9, 10], and few transducer protocols focus on heat extraction from quantum systems to achieve ground state cooling or state preparation. The development of efficient polarization methods is particularly critical for spin ensembles, as conventional techniques often rely on resonant energy exchange methods, such as the Hartman-Hahn transfer [17, 18]. Polarization based on energy transfer typically exploits the interaction between a single spin and all the spins of the ensemble. This single spin can be polarized repeatedly, thus realizing a continuous removal of entropy until the ensemble is finally polarized. The challenge is that if the ensemble spins have similar g-factors, the risk of them being trapped in dark states due to Hamiltonian symmetries is high, preventing complete polarization of the ensemble [18, 19].
In this work, we consider a tripartite hybrid system where an oscillator serves as a transducer for a single probe spin and a spin ensemble. By exploiting the measurement feedback from the frequently measured probe spin, efficient heat extraction from the ensemble can be realized, requiring only weak dispersive (off-resonant) coupling between the spins and the oscillator. The probe control sequence is shown schematically in Fig. 1(b). Achieving complete ensemble polarization by such probe projections can avoid the dark state problem associated with resonant energy transfer. Furthermore, the ability to operate under dispersive interaction is important for various systems whose original dynamics do not allow direct energy exchange [20, 21]. Finally, we show that the same control sequence on the probe can also be used to prepare macroscopic oscillators (e.g., NMO) into specific quantum states that can potentially be used for sensing and error correction applications.
The oscillator here serves as a simplified conceptual model for a common bus (transducer) connecting two distant quantum systems that are not directly or weakly coupled. One can consider a general scenario where the bus is a large spin network that can be effectively approximated as a bosonic system via the Holstein-Primakoff transformation, such as the spin wave excitations (magnons) [22, 11]. This is relevant when a single spin, such as nitrogen-vacancy (NV) centers in diamond [23, 24], is used to polarize or probe distant target spins beyond the distance criteria set by the sensing volume [25]. The target may be distant carbon nuclear spins within the diamond substrate [26, 27], or spins attached to an external protein/molecule [28]. In such cases, probe-target interactions must be mediated by a transducer, such as the spin network schematically shown in the inset of Fig. 1(a) [26, 27, 28, 10]. Despite tuning both the network-to-probe and network-to-target interactions to resonance, the significant inhomogeneities typically present within a spin network make it difficult to become polarized, rendering resonant probe-target energy transfer ineffective [28, 26, 27]. Here, we instantiate such a scenario by considering the oscillator as the fundamental mode of an NMO, specifically the vibration of a clamped cantilever, as schematically shown in Fig. 1(a) [20, 21].
@@
The paper is organized as follows: In Sec. II, we outline the Hamiltonian of the considered hybrid system and discuss possible experimental implementations of the proposed protocol. In Sec. A, we describe the dynamics of the hybrid system and analyze how periodic inversion pulses on the probe spin can significantly modify the dynamics of the ensemble and the oscillator. In Sec. IV, we detail how repeated probe projections can bring the oscillator and the ensemble into their ground states. In particular, we study the parameters that allow cooling of the spin ensemble, the oscillator, or both. Additionally, we discuss how it is possible to prepare the oscillator into a complex quantum state using the same probe control sequence as the cooling protocol. Finally, Sec. V provides a summary and discusses possible directions for future research.
II System Hamiltonian and experimental realization
@@
For simplicity, we assume (though not necessary) that the ensemble spins interact with the common bosonic mode (oscillator) uniformly, with a coupling strength denoted as $g$. In addition, the probe spin is assumed to have a distinct coupling strength with the oscillator, denoted as $g_{0}$. The hybrid system’s Hamiltonian is then formulated as follows:
@@
Figure 1: Schematic representation of the experimental setup and control strategy. (a) Shown is a clamped nano/micromechanical cantilever carrying a single defect spin, which in this study is exemplified by an NV electronic spin. The interaction between the NV center spin and the cantilever motion is mediated by strain, while single qubit gates are achieved by microwave (MW) pulses. Initialization and readout processes are facilitated by optical pulses, with a photon detector used for readout. A magnetic tip at the edge of the cantilever creates a gradient field that allows interaction with a nearby spin ensemble, coupling it to the mechanical motion of the cantilever. In the inset, we schematically represent a network of interacting spins as a quantum bus. When the number of spins in the network is large, it can be modeled as a bosonic system (e.g., magnons) by the Holstein-Primakoff transformation. Note that it is simplified here as the fundamental mode of the cantilever’s vibration. (b) The control sequence of the probe spin to perform the cooling protocol. First, the probe spin is reset to the |$\ket{0}$ state by optical pulses, then a $\pi/2$ MW pulse brings it to a state of equal superposition, which is followed by $n$ inversion MW pulses applied periodically with a carefully chosen time interval $\tau$. Another $\pi/2$ MW pulse is applied before the probe is optically read into the computational base. The above process is repeated $M$ times, bringing the oscillator and the spin ensemble progressively closer to their thermal ground states. Note that the probe to be read must be post-selected in |$\ket{0}$ each time, otherwise the whole process starts over.
This hybrid system can be realized as shown in Fig. 1(a) by a unilaterally clamped cantilever, embedding a single NV center positioned at the clamping point such that it experiences maximum strain and thus higher coupling strength. On the free side of the cantilever, we attach a magnetic tip to generate a strong field gradient that couples to a nearby spin ensemble on an external substrate [9, 6]. In this way, we exploit both possibilities to achieve spin-mechanical coupling [7, 29, 30] by local (strain) and non-local (magnetic-tip) coupling of spins to mechanical modes within a single setup. This allows the distinct manipulation of quantum systems that are otherwise not controllable.
@@
The probe spin is chosen as the electronic spin associated with a NV center due to its good spin coherence properties, high degree of spin control, and well resolved optical spectra at low temperatures [31, 32, 33]. Its optical interface allows excitation to different optically excited energy levels that can either initialize or read out the spin state [34]. The NV electronic spins are also sensitive to their environment, allowing them to be (weakly) coupled to external spins, photons [20, 35] and phonons [21, 16].
@@
The NV electronic spin is spin-1, which can be operated as a qubit with its $m_{s}=0,-1$ sublevels. This requires an external magnetic field aligned with its symmetry axis to separate the $m_{s}=\pm 1$ levels, which is typically around 400 Gauss [32]. Such an arrangement negates the need to consider transverse spin-strain coupling due to its direct interaction between the $m_{s}=\pm 1$ levels [7, 29, 30], allowing us to safely ignore the $m_{s}=+1$ sublevel in this work. What remains is the longitudinal strain coupling acting on the qubit levels, with a number that can be up to 10 kHz for NV centers [36]. This number can be even higher for other solid-state defect centers [37].
@@
In contrast, the spin ensemble to be cooled can be chosen more flexibly, since precise control of these spins is not a requirement for the current study. Potential targets include distant spins on other coupled cantilevers, spins at the clamped edge that are strain-coupled to the mechanical mode, or surface electron spins that are not well controlled. For example, if the ensemble spins are dark spins on the surface of the NMO [38] or other ferromagnetic particles [39], optical manipulation over them can be difficult, highlighting the need for other indirect polarization techniques. Note that this ensemble-oscillator interaction generated by the attached magnetic tip can be (1-2 orders of magnitude) stronger than the spin-strain coupling between the probe and the oscillator [40].
It is important to note that although we specifically consider a cantilever system, an alternative experimental realization could use the spin wave excitations of the spins (magnons) as the transducer [10, 22, 11]. An advantage of this alternative is that the spin-magnon coupling can easily reach a few MHz [41], much stronger than the typical strength of spin-cantilever coupling by either strain or magnetic tip.
III System dynamics and its tuning
A Original system dynamics
@@
Since the ensemble spins form a finite-dimensional Hilbert space, it is convenient to write the system evolution operators in the form [42, 43]:
@@
where $\mathcal{T}$ the time ordering operator; the terms $h_{\scriptscriptstyle k,\pm}(t)$ are the oscillator Hamiltonians conditioning on the spin states, which are expressed as:
$h_{\scriptscriptstyle k,\pm}(t)=\frac{g(k-N/2)\pm g_{0}}{2}\left(\hat{a}e^{-i\omega t}+\hat{a}^{\dagger}e^{i\omega t}\right);$ (4)
and the operator $\mathrm{I}_{\scriptscriptstyle k}$ projects the spin ensemble into the subspace where $k$ spins are pointing up. For example, when $N=3$ and $k=2$, the projector is written as
@@
where $\alpha(t)=1-e^{i\omega t}$ and $\theta_{\scriptscriptstyle k}$ is given by
$\begin{split}\theta_{\scriptscriptstyle k}&=-\frac{\omega t-sin(\omega t)}{\omega t}\frac{\left[g(k-N/2)\pm g_{0}\right]^{2}}{4\omega}.\end{split}$ (7)
B Inversion pulses on probe
In addition to the natural dynamics induced by the Hamiltonian given in Eq. (1), we also apply a periodic sequence of $\pi$ pulses to the probe spin. These pulses significantly alter the dynamics of the system by imposing a time-periodic function:
@@ B Inversion pulses on probe
where $j=0,1,2,\cdots,n$, and $\tau$ denotes the interval between successive pulses. By replacing $g_{0}$ with $f(t)g_{0}$ in $h_{\scriptscriptstyle k,\pm}(t)$, the conditional evolution operators in Eq. (6) are now written as
$\begin{split}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)=&\exp\Big{[}\frac{g(k-N/2)}{2\omega}\left(\alpha(t)a^{\dagger}-\alpha^{*}(t)a\right)\\ &\;\pm\frac{g_{0}}{2\omega}\left(F(\epsilon,t)a^{\dagger}-F^{*}(\epsilon,t)a\right)\Big{]}\times\exp\left[i\theta_{\scriptscriptstyle k}(t)\right],\end{split}$ (9)
@@
$\begin{split}F(t)&=i\omega\ \sum_{k=0}^{n}\int\limits_{k\tau}^{(k+1)\tau}\mathrm{d}t\;(-1)^{k}e^{-i\omega t}\\ &=\left[1-(-1)^{n+1}e^{-i(n+1)\omega\tau}+2\sum_{k=1}^{n}(-1)^{k}e^{-ik\omega\tau}\right].\end{split}$ (10)
This filter function is centered at $\tau=\pi/\omega$ with its width decreasing approximately as $1/n^{2}$ [43, 42, 44] and asymptotically converges to a delta function:
$\lim\limits_{n\rightarrow\infty}F(t)=2(n+1)\delta(\omega-\frac{\pi}{\tau}),$ (11)
which supports our initial assumption that all other phonon modes are effectively decoupled in Eq. (1) with large number of pulses $n$. In addition, we also want the maximum possible coupling strength between them. One can see from Eq. (9) that $|g_{0}F(t)|$ should be comparable to the oscillator frequency $\omega$, which requires at least $n\approx\mathcal{O}(\omega/g_{0}$) [43].
To achieve some tunability of the system dynamics that facilitates our proposed cooling protocol, we introduce the pulse detuning $\epsilon$ as a control parameter:
$\tau=\frac{\pi}{\omega-\epsilon},\text{ with }\frac{\epsilon}{\omega}\ll 1.$ (12)
Note that the detuning is intentionally kept significantly smaller than the oscillator frequency, so that the inversion pulses are in the near-resonance regime. Accordingly, the filter function can now be simplified as
@@
For higher values of pulse detuning $\epsilon$, making the above equation closer to a delta function would require an even larger pulse number $n$.
Figure 2: (a) The quantum circuit progressively projects the oscillator ensemble system by projection of the probe spin. The pulse sequence on the physical level is shown in fig. 1(b). In each repetition, the probe spin is first prepared in the equal superposition state $\ket{+}$ by the optical reset and a microwave Hadamard gate. Then repeated inversion pulses are applied, resulting in a controlled oscillator displacement $\sum_{k}\mathcal{D}^{f}_{\scriptscriptstyle k,\pm}(t)\otimes\mathrm{I}_{k}$. Here the subscript $\pm$ depends on the spin state of the probe and $k$ on the number of upward-pointing ensemble spins. The procedure ends with a Hadamard gate and a measurement of the probe spin, with the result set to 0. Such a repetition effectively implements the projector given in Eq. (14). (b) An example of how an ensemble of $N=14$ spins is polarized after $M=100$ probe projections. The spins are initially in a completely mixed state, which is projected to the state with all spins pointing up with a probability of 95.82%. Note that achieving such a high degree of polarization requires a carefully chosen pulse detuning, which in this case is $\epsilon=1.16$ kHz. Further simulation details are summarized in Appendix A.
C Tuning dynamics through probe projections
Although we focus specifically on the weak dispersive coupling between each spin and the oscillator, we deliberately ignore any direct interactions between the spins, making energy transfer between the probe and the ensemble unlikely. Nevertheless, by carefully choosing the pulse detuning $\epsilon$ and the number of pulses $n$, energy transfer-like changes in the dynamics of the spin ensemble can be induced, as we will discuss below.
The dynamics modification is facilitated by a cyclic process, involving the projective measurement and reinitialization of the probe spin, as schematically shown in Fig. 1(a). First, the probe is prepared in the state $\ket{+}$. Then the system undergoes an evolution for a time $t=n\tau$ under repeated inversion pulses, as described in Eq (3). Finally, the probe spin is measured in the computational basis after an additional $\pi/2$ pulse is applied to it.
The process described above is represented by the circuit in Fig. 2(a), effectively implementing oscillator displacements conditioning on the probe state. The whole circuit is then equivalent to the implementation of a projector on the oscillator-ensemble system, which is mathematically formulated as:
$\mathcal{P}_{\scriptscriptstyle N,\pm}=\sum_{\scriptscriptstyle k=0}^{\scriptscriptstyle N}V_{k,\pm}\otimes\mathrm{I}_{k},\quad V_{k,\pm}=\frac{1}{2}\left(\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,+}(t)\pm\mathcal{D}^{\scriptscriptstyle f}_{\scriptscriptstyle k,-}(t)\right).$ (14)
The operator $V_{k,\pm}$ is a projector on the oscillator, where $k$ is the number of ensemble spins pointing up, and the subscript $\pm$ depends on the measurement result of the probe spin.
Note that such repeated projections of the probe is central to the oscillator cooling protocol in our previous work [43], which is now extended to cooling/polarizing a spin ensemble. Here, we start with $N$ ensemble spins initially in a fully mixed state, and the oscillator initially in its thermal state. These initial states are written as